Suppose your project is worth $822 with a probability of 0.9 and $492 with a probability of 1-0.9. The appropriate cost of capital is 9.2% for the overall project. If you want to raise $664 today and promise to bond investors a rate of return of $5.4, what's the expected rate of return of this bond?____________________ Please answer both Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1
A B C D E F G H I
2
3 Bond holders are paid the full amount if project worth is higher than the total debt value,
4 however if the value of the project is lower than the value of the project then the
5 debt value equals the project value.
6
7 Amount of Debt Raised $664
8 Promised return to debtholders 5.40%
9 Expected Value of Debt after one Year $699.86 =D7*(1+D8)
10
11 Probability Project Worth Debt Value Return to debt holders
12 0.9 $822 $699.86 5.40% =(E12-$D$7)/$D$7
13 0.1 $492 $492.00 -25.90% =(E13-$D$7)/$D$7
14
15 Expected return of debtholder =0.9*5.40%+0.1*-25.90%
16 2.27% =C12*F12+C13*F13
17
18 Expected return of debtholder 2.27%
19